1.An airline records flight delays in and out of Chicago O'Hare airport for a year. The average delay for these flights is 9.05 minutes with a standard deviation of 4.04 minutes. For a sample of 61 flights, 23% of flights will have an average delay less than how many minutes?

a.There is not enough information

b. 12.03

c. 6.07

d.9.43

e.8.67

2.As a result of the subprime mortgage crisis of 2008, approximately 11.81% of new mortgage originations went into foreclosure by 2009. Suppose in the Kalamazoo area there were 2,587 homes in which the mortgage originated during this period. What is the probability that less than 12.37% of the homes would be foreclosed on?

a.0.1887

b.>0.999

c.0.8113

d. 42.5921

e.0.5000
Answer #1

Solution:

Question 1)

Given:

Mean = \mu =9.05

Standard Deviation = \sigma = 4.04

Sample size= n= 61

Find sample mean \bar{x} such that:

P(\bar{X}<\bar{x}) = 0.23

Thus find z value such that:

P( Z< z) =0.23

Look in z  table for Area = 0.2300 or its closest area and find corresponding z value.

Area -0.2296 is closest to 0.2300 and it corresponds to -0.7 and 0.04

thus z = -0.74

Now use following formula to find \bar{x } value:

\bar{x } = \mu + z \times \sigma /\sqrt{n}

\bar{x } = 9.05 + (-0.74) \times 4.04 /\sqrt{61}

\bar{x }= 9.05 + (-0.74) \times 4.04 /7.81025

\bar{x } = 9.05 - 0.38

\mathbf{{\color{DarkOrange} \bar{x }=8.67}}

Question 2)

Given:

p = 0.1181

n = 2587

Find:

P(\hat{p}<0.1237)=...........?

Find z score for 0.1237

z= \frac{\hat{p}-p}{\sqrt{\frac{p\times (1-p)}{n}}}

z= \frac{0.1237-0.1181}{\sqrt{\frac{0.1181 \times (1-0.1181)}{2587}}}

z= \frac{0.0056 }{\sqrt{\frac{0.1181 \times 0.8819 }{2587}}}

z= \frac{0.0056 }{\sqrt{ 0.00004026 }}

z= \frac{0.0056 }{0.0063451 }

z= 0.882575

z= 0.88

Thus we get:
P(\hat{p}<0.1237)=P( Z < 0.88)

Look in z table for z =0.8 and 0.08 and find corresponding area.

.00 .01 .02 .03 .04 .05 .06 .07 .08 .09 Z $319 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 5000 5040 5080 .5120 5160 5199 5239 .5279

P( Z < 0.88) = 0.8106

closest answer is: 0.8113 thus correct answer is: 0.8113

Excel command gives accurate answer:

=NORM.S.DIST( z ,cumulative)

=NORM.S.DIST(0.882575,TRUE)

=0.8113